# The Catalogue of Virtual Early-Type Galaxies from IllustrisTNG: Validation and Real Observation Consistency

### v1

@@ 3.2.1 Definitions
$\gamma(r)=-\frac{\partial\log\rho(r)}{\partial\log r},$ (23)
@@
$\gamma^{\rm MW}(r)=3-\frac{\partial\log M(r)}{\partial\log r}.$ (25)
@@ 3.2.2 Results and comparison with literature
In Table 2 we report some statistical properties of the slopes for the total mass density distributions and also the three components, i.e., star, gas and dark matter, separately. We also report the $\gamma_{\rm tot,2}^{\rm PL}$ values discussed in §2.4 and the average slope defined in Eq. (21): these latter estimates are, by construction, better references for the observed values as they are defined on a radial sample closer to the typical values adopted for real systems. In the same table, we include the mean value, the stellar mass-weighted (SMW for short) mean – where the weights assigned for each virtual-ETG system are their 2D total stellar mass $M_{2D,\star}^{\rm tot}$ –, median value, standard error (Std. Err.) of the mean and the observed scatter. The distributions for the stellar, dark matter and total slopes are shown in Fig. 9. These are the most relevant quantities we will discuss in the following, having already mentioned that the $\gamma_{\rm gas}$ is little robust and poorly physically meaningful for ETG samples. Looking at the distribution of our results in Fig. 9 and the statistics in Table 2, we find that the $\gamma_{\rm tot}^{\rm PL}$ is nicely “isothermal”, according to the different statistical indicators, while $\gamma_{\rm 2,tot}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$ are slightly steeper than 2, although in all the cases “isothermality” is included in the large scatter (0.148, 0.203 and 0.195 for $\gamma_{\rm tot}$, $\gamma_{\rm tot,2}$ and $\gamma_{\rm tot}^{\rm AV}$, respectively). The reason for $\gamma_{\rm tot}^{\rm PL}$ to be shallower than $\gamma_{\rm tot,2}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$ comes from the fact that the last two slopes are defined over a radial range where the steeper stellar density profile ($\gamma_{\star}\sim 3$, see Table 2) dominates over the DM density profile which has a shallower slope ($\gamma_{\rm DM}\sim 1.7$, see also Table 2 ).
Figure 9: Distribution of the slopes for the stellar (shades of orange), dark matter (shades of blue) and total components (shades of red). The power-law (PL) slopes are computed according to Eq. (14) over the radial range $[0.4,2]r_{1/2}$ for $\gamma_{i}^{\rm PL}$ and $[0.4,4]r_{1/2}$ for $\gamma_{i,2}^{\rm PL}$. The average (AV) slopes are computed according to Eq. (21) over the radial range $[0.4,2]r_{1/2}$.
@@
Dark Matter. For the DM component, we have found (mean $\pm$ scatter) $\langle\gamma_{\rm DM}^{\rm PL}\rangle\pm\sigma_{\rm DM}=1.742\pm 0.103$, $\langle\gamma_{\rm DM,2}^{\rm PL}\rangle\pm\sigma_{\rm DM,2}=1.717\pm 0.135$ and $\langle\gamma_{\rm DM}^{\rm AV}\rangle\pm\sigma_{\rm DM}=1.714\pm 0.163$. This component is the one with the minimum slope scatter in each respective definition. The DM also has a shallower radial distribution compared to the stellar and total slopes (see Table 2). Using TNG100 ETGs at $z=0$, and considering the radial range $0.4r_{1/2}\leq r\leq 4r_{1/2}$, Wang et al. (2020) found $\langle\gamma_{\rm DM}^{\rm PL}\rangle=1.760\pm 0.005$ and $\sigma_{\rm DM}=0.108$ for early-type galaxies with stellar masses in the range $10^{10.7}\leq M_{\star}/M_{\odot}\leq 10^{11.9}$, which is unsurprisingly consistent with the one found in our work , as in our analysis we are closely following their procedure, and just differing for the range of stellar masses adopted that in our case includes ETGs with masses as small as $10^{10.3}M_{\odot}$ and up to $10^{11.9}M_{\odot}$. With Illustris-1, using different radial and redshift intervals (again, $0.5R_{\rm e}<r<2R_{\rm e}$ and $z=0.3$), Xu et al. (2017) found $\langle\gamma_{\rm DM}^{\rm PL}\rangle=1.58$ with a scatter $\sigma_{\rm DM}=0.17$, i.e., shallower than the results found in TNG100, but yet in agreement within $1\sigma$, as we illustrate in Fig. 10. The shallow value obtained from Illustris-1 is likely because $\gamma_{\rm DM}$ steepens by increasing the radial interval where the density profile is defined, as suggested by the results in Wang et al. (2020). Thus, for a dark matter density slope defined within $R_{\rm e}$, one should expect a shallower $\gamma_{\rm DM}$. The results of TNG100 above suggest a contraction of the dark matter halo from a “collisionless” NFW-like DM slope $\sim-1$, in agreement with the results of Napolitano et al. (2010, see discussion in their §3.2) (see also Tortora et al. 2010 for further discussion) which, assuming a Chabrier IMF, found an average $\langle\gamma_{\rm DM}\rangle\sim 1.6\pm 0.2$ from the Jeans dynamical analysis of a local sample of ETGs ($z\lesssim 0.03$). A consistent result has also been found by Sonnenfeld et al. (2012) who measured a dark matter slope $\gamma_{\rm DM}^{\rm AV}=1.7\pm 0.2$ in a strong lensed ETG at $z=0.222$. From strong lensing again, using the SLACS sample (Bolton et al., 2006) at $z<0.3$, and assuming density distribution inside the effective radius, Grillo (2012) have found a mean dark matter slope of $\langle\gamma_{\rm DM}^{\rm AV}\rangle=1.7\pm 0.5$, in agreement with our estimate and Sonnenfeld et al. (2012). However, Dutton & Treu (2014) found a contradicting result using the same Grillo (2012) sample and claimed a mean $\langle\gamma_{\rm DM}^{\rm MW}\rangle=1.40^{+0.15}_{-0.26}$ interior to $R_{\rm e}$. The main difference between the Dutton et al. analysis and the one mentioned above resides in the different IMF adopted, with the latter assuming a Salpeter IMF, which favours mild contraction. This is also found by Newman et al. (2015) which, using a Salpeter IMF and combining data from 10 group lenses, found $\langle\gamma_{\rm DM}^{\rm MW}\rangle=1.35\pm 0.09$ inside $R_{\rm e}$ for ETGs within $10^{13}-10^{15}M_{\odot}$ halos, mean redshift $\langle z\rangle=0.36$ , also favouring mild contraction. This IMF vs. DM degeneracy is a well known issue in both dynamical (e.g., Napolitano et al. 2010) and strong lensing analysis (e.g., Treu et al. 2010), which also requires uniformity of approaches/assumptions when it comes to comparison among different analyses. In this respect, our result from TNG100 allows us to set a reference for the inferences based on Chabrier IMF.
@@
Total mass. Moving to the total mass density slope, we have found that (mean $\pm$ scatter) $\langle\gamma_{\rm tot}^{\rm PL}\rangle=1.998\pm 0.148$, $\langle\gamma_{\rm tot,2}^{\rm PL}\rangle=2.055\pm 0.203$ and $\langle\gamma_{\rm tot}^{\rm AV}\rangle=2.057\pm 0.195$. These results show a nice agreement with a singular isothermal sphere model ($\gamma=2$ in Eq. 14). In each definition, the slightly larger scatter of $\gamma_{\rm tot}$ with respect to $\gamma_{\rm DM}$ and $\gamma_{\star}$ is likely inherited from the gas distribution, which is the greatest source of stochasticity among the three components. Using the same radial range ($0.4r_{1/2}<r<4r_{1/2}$), Wang et al. (2020) found $\langle\gamma_{\rm tot}^{\rm PL}\rangle=2.011$ and $\sigma_{\rm tot}=0.171$, which (again) is in perfect agreement with our result. Using Illustris-1, Xu et al. (2017) found a stellar mass-weighted mean total slope $\langle\gamma_{\rm tot}^{\rm PL}\rangle_{\rm SMW}=2.07$ and scatter $\sigma_{\rm tot}=0.26$, again, for galaxies at $z=0.3$ and over the radial range $0.5R_{\rm e}<r<2R_{\rm e}$, which is steeper compared to ours ($\langle\gamma_{\rm tot}\rangle_{\rm SMW}=1.939$, see Table 2). As mentioned, we have also estimated the mean total slopes get by using $2r_{1/2}$ as the outer radial limit, namely $\gamma_{\rm tot,2}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$, and found $\langle\gamma_{\rm tot,2}^{\rm PL}\rangle=2.055\pm 0.203$, and $\langle\gamma_{\rm tot}^{\rm AV}\rangle=2.057\pm 0.195$. These are slightly steeper than the result considering an upper bound of $4r_{1/2}$, as also seen in Fig. 10. This is expected as the shorter radial range makes the total slope more aligned to the steeper stellar density slope, dominating in the very centre of galaxies.
@@ 3.3 Total density slope correlations
In this section, we quantitatively compare the correlations of the total slopes and the other galaxy parameters derived for the virtual-ETG sample with the same relations derived from observational data. This is a very stringent validation test for our catalogue, as we use a full set of observational-like parameters to derive scaling relations from simulated galaxies. In particular, we compute the correlation of the total density slope with some Sérsic parameters (i.e., $n,R_{\rm e}$ and the derived stellar mass $M_{2D,\star}^{\rm tot}$) and kinematics, i.e., $\sigma_{\rm e}$, using the different slope definition adopted as in §3.2.1. In the analysis, we denote the slopes of the best-fit line as $\partial\gamma_{\rm tot}/\partial{X}$, for a particular parameter $X$ and the strength of the correlation by the Pearson’s correlation coefficient $r$ [^12] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.pearsonr.html. We regard a correlation as being significant if the p-value of the best-fit linear regression is less than or equal to $0.01$ (see, e.g., Moore et al. 2009 for a statistical framework). This ensures that we can reject the null hypothesis, which, in our case, states that there is no correlation in the data.
@@
For the SPIDER sample we use the Sérsic index, half-light radii, stellar mass (assuming a Chabrier IMF), stellar velocity dispersion within $R_{\rm e}$ and mass-weighted density slopes within $R_{\rm e}$ from Tortora et al. (2014), in which ETGS have stellar masses $\gtrsim 10^{10}~{}M_{\odot}$ For the SL2S sample, we use the logarithmic density slopes, stellar masses assuming a Salpeter IMF and effective radius from Sonnenfeld et al. (2013) for objects with stellar masses $\gtrsim 10^{10.8}~{}M_{\odot}$ for $0.80.2<z<$ 0.8. For consistency, we converted these stellar masses to a Chabrier using $M_{\star,\text{Chab}}=0.61M_{\star,\text{Salp}}$ (Madau & Dickinson, 2014), where $M_{\star,\text{Chab}}$ is the stellar mass derived assuming a Chabrier IMF and $M_{\star,\text{Salp}}$ is the one derived assuming a Salpeter IMF. For SLACS sample, we extract the power-law slope, effective radius, and stellar mass with Chabrier IMF from Auger et al. (2010), where the ETGs have stellar masses $\gtrsim 10^{10}M_{\odot}$ and the redshift range is $0.780.24<z<$ 0.78. Although these ETGs samples have (projected) total stellar masses approximately equivalent to ours, we selected only those with $M_{\rm 2D,\star}^{\rm tot}\geq 10.5$ to ensure a more consistent comparison. Moreover, for Tortora et al. (2014), we selected only those galaxies with $n\geq 2$ and for Zhu et al. (2023b) those with $2\leq n\leq 5.9$ (see §3.1). After filtering the relevant data from the samples, we are left with $504$ objects for Zhu et al. (2023b), $16$ for Serra et al. (2016), $25$ for Sonnenfeld et al. (2013), $59$ for Auger et al. (2010), and the greatest dataset is from Tortora et al. (2014), with $4112$ galaxies.
@@ 3.3.2 Correlation with effective radius
In general, the direction of the $\gamma-R_{\rm e}$ correlation is consistent across all samples. However, the DynPop (Zhu et al., 2023b) sample with a free NFW and a gNFW model for the dark matter distribution represents the weakest correlations, which are, in fact, negligible, having a p-value below the threshold we imposed. On the other hand, the ETGs from DynPop considering a fixed NFW model show consistency with the virtual-ETG, although shallower, with a slope of $\partial\gamma_{\rm tot}^{\rm MW}/\partial\log R_{\rm e}=-0.556\pm 0.019$ and a Pearson’s correlation coefficient of $-0.788$. This slope is consistent with our result for the correlation between the effective radius and the slope $\gamma_{\rm tot}^{\rm PL}$ at the $1\sigma$ confidence level.
@@ 3.4 Projected versus three-dimensional quantities
In this section, we focus on the study of the correlations between the three-dimensional quantities and their projected counterparts or equivalent. In particular, we analyse the relation between the effective radius and the stellar half-light radius (which we approximate as being the stellar half-mass radius $r_{1/2}$), the relation between central Sérsic luminosity and FSPS luminosity and also between central projected and 3D stellar mass. For the luminosity and stellar mass, the central values are defined within apertures of $R_{\rm e}$ and $2R_{\rm e}$.

### v2

@@ 3.2.1 Definitions
$\gamma(r)=-\frac{\rm d\log\rho(r)}{\rm d\log r},$ (23)
@@
$\gamma^{\rm MW}(r)=3-\frac{\rm d\log M(r)}{\rm d\log r}.$ (25)
@@ 3.2.2 Results and comparison with literature
In Table 2 we report some statistical properties of the slopes for the total mass density distributions and also the three components, i.e., star, gas and dark matter, separately. We also report the $\gamma_{\rm tot,2}^{\rm PL}$ values discussed in §2.4 and the average slope defined in Eq. (21): these latter estimates are, by construction, better references for the observed values as they are defined on a radial sample closer to the typical values adopted for real systems. In the same table, we include the mean value, the stellar mass-weighted (SMW for short) mean – where the weights assigned for each virtual-ETG system are their 2D total stellar mass $M_{2D,\star}^{\rm tot}$ –, median value, standard error (Std. Err.) of the mean and the observed scatter. These are the most relevant quantities we will discuss in the following, having already mentioned that the $\gamma_{\rm gas}$ is little robust and poorly physically meaningful for ETG samples. Looking at the statistics in Table 2, we find that the $\gamma_{\rm tot}^{\rm PL}$ is nicely “isothermal”, according to the different statistical indicators, while $\gamma_{\rm 2,tot}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$ are slightly steeper than 2, although in all the cases “isothermality” is included in the large scatter (0.148, 0.203 and 0.195 for $\gamma_{\rm tot}$, $\gamma_{\rm tot,2}$ and $\gamma_{\rm tot}^{\rm AV}$, respectively). The reason for $\gamma_{\rm tot}^{\rm PL}$ to be shallower than $\gamma_{\rm tot,2}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$ comes from the fact that the last two slopes are defined over a radial range where the steeper stellar density profile ($\gamma_{\star}\sim 3$, see Table 2) dominates over the DM density profile which has a shallower slope ($\gamma_{\rm DM}\sim 1.7$, see also Table 2 ).
@@
Dark Matter. For the DM component, we have found (mean $\pm$ scatter) $\langle\gamma_{\rm DM}^{\rm PL}\rangle\pm\sigma_{\rm DM}=1.742\pm 0.103$, $\langle\gamma_{\rm DM,2}^{\rm PL}\rangle\pm\sigma_{\rm DM,2}=1.717\pm 0.135$ and $\langle\gamma_{\rm DM}^{\rm AV}\rangle\pm\sigma_{\rm DM}=1.714\pm 0.163$. This component is the one with the minimum slope scatter in each respective definition. The DM also has a shallower radial distribution compared to the stellar and total slopes (see Table 2). Using TNG100 ETGs at $z=0$, and considering the radial range $0.4r_{1/2}\leq r\leq 4r_{1/2}$, Wang et al. (2020) found $\langle\gamma_{\rm DM}^{\rm PL}\rangle=1.760\pm 0.005$ and $\sigma_{\rm DM}=0.108$ for early-type galaxies with stellar masses in the range $10^{10.7}\leq M_{\star}/M_{\odot}\leq 10^{11.9}$, which is unsurprisingly consistent with the one found in our work , as in our analysis we are closely following their procedure, and just differing for the range of stellar masses adopted that in our case includes ETGs with masses as small as $10^{10.3}M_{\odot}$ and up to $10^{11.9}M_{\odot}$. With Illustris-1, using different radial and redshift intervals (again, $0.5R_{\rm e}<r<2R_{\rm e}$ and $z=0.3$), Xu et al. (2017) found $\langle\gamma_{\rm DM}^{\rm PL}\rangle=1.58$ with a scatter $\sigma_{\rm DM}=0.17$, i.e., shallower than the results found in TNG100, but yet in agreement within $1\sigma$, as we illustrate in Fig. 9. The shallow value obtained from Illustris-1 is likely because $\gamma_{\rm DM}$ steepens by increasing the radial interval where the density profile is defined, as suggested by the results in Wang et al. (2020). Thus, for a dark matter density slope defined within $R_{\rm e}$, one should expect a shallower $\gamma_{\rm DM}$. The results of TNG100 above suggest a contraction of the dark matter halo from a “collisionless” NFW-like DM slope $\sim-1$, in agreement with the results of Napolitano et al. (2010, see discussion in their §3.2) (see also Tortora et al. 2010 for further discussion) which, assuming a Chabrier IMF, found an average $\langle\gamma_{\rm DM}\rangle\sim 1.6\pm 0.2$ from the Jeans dynamical analysis of a local sample of ETGs ($z\lesssim 0.03$). A consistent result has also been found by Sonnenfeld et al. (2012) who measured a dark matter slope $\gamma_{\rm DM}^{\rm AV}=1.7\pm 0.2$ in a strong lensed ETG at $z=0.222$. From strong lensing again, using the SLACS sample (Bolton et al., 2006) at $z<0.3$, and assuming density distribution inside the effective radius, Grillo (2012) have found a mean dark matter slope of $\langle\gamma_{\rm DM}^{\rm AV}\rangle=1.7\pm 0.5$, in agreement with our estimate and Sonnenfeld et al. (2012). However, Dutton & Treu (2014) found a contradicting result using the same Grillo (2012) sample and claimed a mean $\langle\gamma_{\rm DM}^{\rm MW}\rangle=1.40^{+0.15}_{-0.26}$ interior to $R_{\rm e}$. The main difference between Dutton & Treu (2014) analysis and the one mentioned above resides in the different IMF adopted, with the latter assuming a Salpeter IMF, which favours mild contraction. This is also found by Newman et al. (2015) which, using a Salpeter IMF and combining data from 10 group lenses, found $\langle\gamma_{\rm DM}^{\rm MW}\rangle=1.35\pm 0.09$ inside $R_{\rm e}$ for ETGs within $10^{13}-10^{15}M_{\odot}$ halos, mean redshift $\langle z\rangle=0.36$ , also favouring mild contraction. This IMF vs. DM degeneracy is a well known issue in both dynamical (e.g., Napolitano et al. 2010) and strong lensing analysis (e.g., Treu et al. 2010), which also requires uniformity of approaches/assumptions when it comes to comparison among different analyses. In this respect, our result from TNG100 allows us to set a reference for the inferences based on Chabrier IMF.
@@
Total mass. Moving to the total mass density slope, we have found that (mean $\pm$ scatter) $\langle\gamma_{\rm tot}^{\rm PL}\rangle=1.998\pm 0.148$, $\langle\gamma_{\rm tot,2}^{\rm PL}\rangle=2.055\pm 0.203$ and $\langle\gamma_{\rm tot}^{\rm AV}\rangle=2.057\pm 0.195$. These results show a nice agreement with a singular isothermal sphere model ($\gamma=2$ in Eq. 14). In each definition, the slightly larger scatter of $\gamma_{\rm tot}$ with respect to $\gamma_{\rm DM}$ and $\gamma_{\star}$ is likely inherited from the gas distribution, which is the greatest source of stochasticity among the three components. Using the same radial range ($0.4r_{1/2}<r<4r_{1/2}$), Wang et al. (2020) found $\langle\gamma_{\rm tot}^{\rm PL}\rangle=2.011$ and $\sigma_{\rm tot}=0.171$, which (again) is in perfect agreement with our result. Using Illustris-1, Xu et al. (2017) found a stellar mass-weighted mean total slope $\langle\gamma_{\rm tot}^{\rm PL}\rangle_{\rm SMW}=2.07$ and scatter $\sigma_{\rm tot}=0.26$, again, for galaxies at $z=0.3$ and over the radial range $0.5R_{\rm e}<r<2R_{\rm e}$, which is steeper compared to ours ($\langle\gamma_{\rm tot}\rangle_{\rm SMW}=1.939$, see Table 2). As mentioned, we have also estimated the mean total slopes by using $2r_{1/2}$ as the outer radial limit, namely $\gamma_{\rm tot,2}^{\rm PL}$ and $\gamma_{\rm tot}^{\rm AV}$, and we found that they are slightly steeper than the result considering an upper bound of $4r_{1/2}$, as also seen in Fig. 9. This is expected as the shorter radial range makes the total slope more aligned to the steeper stellar density slope, dominating in the very centre of galaxies.
@@ 3.3 Total density slope correlations
In this section, we quantitatively compare the correlations of the total slopes and the other galaxy parameters derived for the virtual-ETG sample with the same relations derived from observational data. This is a very stringent validation test for our catalogue, as we use a full set of observational-like parameters to derive scaling relations from simulated galaxies. In particular, we compute the correlation of the total density slope with some Sérsic parameters (i.e., $n,R_{\rm e}$ and the derived stellar mass $M_{2D,\star}^{\rm tot}$) and kinematics, i.e., $\sigma_{\rm e}$, using the different slope definitions adopted as in §3.2.1. In the analysis, we denote the slopes of the best-fit line as $\partial\gamma_{\rm tot}/\partial{X}$, for a particular parameter $X$ and the strength of the correlation by the Pearson’s correlation coefficient $r$ [^13] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.pearsonr.html. We regard a correlation as being significant if the p-value of the best-fit linear regression is less than or equal to $0.01$ (see, e.g., Moore et al. 2009 for a statistical framework). This ensures that we can reject the null hypothesis, which, in our case, states that there is no correlation in the data.
@@
For the SPIDER sample we use the Sérsic index, half-light radii, stellar mass (assuming a Chabrier IMF), stellar velocity dispersion within $R_{\rm e}$ and mass-weighted density slopes within $R_{\rm e}$ from Tortora et al. (2014), in which ETGS have stellar masses $\gtrsim 10^{10}~{}M_{\odot}$ For the SL2S sample, we use the logarithmic density slopes, stellar masses assuming a Salpeter IMF and effective radius from Sonnenfeld et al. (2013) for objects with stellar masses $\gtrsim 10^{10.8}~{}M_{\odot}$ for $0.80.2<z<$ 0.8. For consistency, we converted these stellar masses to a Chabrier using $M_{\star,\text{Chab}}=0.61M_{\star,\text{Salp}}$ (Madau & Dickinson, 2014), where $M_{\star,\text{Chab}}$ is the stellar mass derived assuming a Chabrier IMF and $M_{\star,\text{Salp}}$ is the one derived assuming a Salpeter IMF. For SLACS sample, we extract the power-law slope, effective radius, and stellar mass with Chabrier IMF from Auger et al. (2010), where the ETGs have stellar masses $\gtrsim 10^{10}M_{\odot}$ and the redshift range is $0.780.24<z<$ 0.78. Although these ETGs samples have (projected) total stellar masses approximately equivalent to ours, we selected only those with $M_{\rm 2D,\star}^{\rm tot}\geq 10.5$ to ensure a more consistent comparison. Moreover, for Tortora et al. (2014), we selected only those galaxies with $n\geq 2$ and for Zhu et al. (2023b) those with $2\leq n\leq 5.9$ (see §3.1). After filtering the relevant data from the samples, we are left with $504$ objects for Zhu et al. (2023b), $25$ for Sonnenfeld et al. (2013), $59$ for Auger et al. (2010), and the greatest dataset is from Tortora et al. (2014), with $4112$ galaxies.
@@ 3.3.2 Correlation with effective radius
In general, the direction of the $\gamma-R_{\rm e}$ correlation is consistent across all samples. However, the DynPop (Zhu et al., 2023b) sample with a free NFW and a gNFW model for the dark matter distribution represents the weakest correlations, which are, in fact, negligible, having a p-value above the threshold we imposed. On the other hand, the ETGs from DynPop considering a fixed NFW model show consistency with the virtual-ETG, although shallower, with a slope of $\partial\gamma_{\rm tot}^{\rm MW}/\partial\log R_{\rm e}=-0.556\pm 0.019$ and a Pearson’s correlation coefficient of $-0.788$. This slope is consistent with our result for the correlation between the effective radius and the slope $\gamma_{\rm tot}^{\rm PL}$ at the $1\sigma$ confidence level.
@@ 3.4 Projected versus three-dimensional quantities
In this section, we focus on the study of the correlations between the three-dimensional quantities and their projected counterparts or equivalent. In particular, we analyse the relation between the effective radius and the stellar half-light radius $r_{1/2}$, the relation between central Sérsic luminosity and FSPS luminosity and also between central projected and 3D stellar mass. For the luminosity and stellar mass, the central values are defined within apertures of $R_{\rm e}$ and $2R_{\rm e}$.